a mason can make a wall in 6 hours. His helper takes 9 hours. In how many hours can they make the wall together.
Answer by checkley75(3666)

6*9/(6+9)=54/15=3.6 HOURS TO COMPLETE THE WALL WORKING TOGETHER
(THIS IS KNOWN AS THE PRODUCT/SUM SOLUTION)